OpenStudy (anonymous):

If the difference between a number and 7 is doubled, the result is 8 more than the number. Find the number,.?? HELP

OpenStudy (chihiroasleaf):

@kjmoney

OpenStudy (aripotta):

"a number" will be referred to as x. 2(x - 7) = 8 + x 2x - 14 = 8 + x 2x = 22 + x x = 22 so, the number is 22.

OpenStudy (aripotta):

do you see how i got the answer?

OpenStudy (anonymous):

YES.:) THANK YOU> SO THEM YOU DONT DIVIDE IT BY 2X?

OpenStudy (aripotta):

no, i subtracted x on both sides

OpenStudy (aripotta):

since it was + x on the right, i did the opposite to get rid of it

OpenStudy (anonymous):

OH OK. I see. thank you soo much. I really appreciate.

OpenStudy (aripotta):

no problem :)

OpenStudy (anonymous):

:)